From a survey of 600 college students, a marketing research company found that 400 students owned stereos, 200 owned cars, and 75 owned both cars and stereos.
(a) How many students owned either a car or a stereo (but not both)?
(b) How many students do not own either a car or a stereo?

Answer by checkley77(12844)

(400+200)-75=600-75=525 students owned a car or a stereo but not both.
600-525=75 students did not own either a car or a stereo.